OpenStudy (anonymous):

Frank kicks a soccer ball off the ground and in the air with an initial velocity of 30 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 14.1 feet 13.7 feet 13.2 feet 15.2 feet

OpenStudy (anonymous):

ok will use calculus based approach

OpenStudy (anonymous):

take derivative h'(t)=-32t+v

OpenStudy (anonymous):

set it equal to zero to get the critical point

OpenStudy (anonymous):

-32t+v=0 but v=30 ft given 32t=v 32t=30 t=30/32

OpenStudy (anonymous):

are you getting this @MeganLewis ??

OpenStudy (anonymous):

yes. but t=30/32 doesnt equal any of my choices.

OpenStudy (anonymous):

well for that wait a little :P

OpenStudy (anonymous):

t=30/32=0.9375

OpenStudy (anonymous):

put that value of t=0.9375 in the original equation H(t) = −16t2 + vt + s and let me know what you got

OpenStudy (anonymous):

and yes use v=30 and s=0 in the above

OpenStudy (anonymous):

H(0.9375)=-16(0.9375)^2+30(0.9375)+0

OpenStudy (anonymous):

H=14.0625 feet

OpenStudy (anonymous):

is it ok??

OpenStudy (anonymous):

14.1 is the close one

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

yw:)